OpenStudy (anonymous):

Some help? Find an explicit rule for the nth term of the sequence. 1, 3, 9, 27...

jimthompson5910 (jim_thompson5910):

fill in the blank the first term is _______

OpenStudy (anonymous):

First term is one, or do you mean term in equation?

jimthompson5910 (jim_thompson5910):

first term in the given sequence above it is 1, so a = 1

jimthompson5910 (jim_thompson5910):

the common ratio is _______

OpenStudy (anonymous):

3.

jimthompson5910 (jim_thompson5910):

so r = 3

jimthompson5910 (jim_thompson5910):

from here, you plug in a = 1 and r = 3 into \[\large a_{n} = a*r^{n-1}\] to get \[\large a_{n} = a*r^{n-1}\] \[\large a_{n} = 1*r^{n-1}\] \[\large a_{n} = 1*3^{n-1}\] which is the general nth term

OpenStudy (anonymous):

its reeally that simple?

jimthompson5910 (jim_thompson5910):

yep